Two infinite parallel planes have uniform charge density σA = σ and σB = 2σ. There is a conducting plate placed a distance s from plane A. The plate has thickness d. (a)Compute the electric field in region II. (b)Compute the electric field in region III. (c)Compute the potential difference, ∆VII , between the left plane and the conductor. (d)Compute the potential difference, ∆VIII , across the conductor.
Answer #1

a) electric field in region II = σ/2ε0 - 2σ/2ε0

                                             = - σ/2ε0   N/C

b)   electric field in region III   = 0 N/C

c)   The potential difference, ∆VII , between the left plane and the conductor = s * (σ/2ε0 )

                                                                                                                          = (s σ)/(2ε0)     V

d)   potential difference, ∆VIII , across the conductor = d * (σ/2ε0 )

                                                                                  =    (d σ)/(2ε0)     V